# 3DGS.zip: A survey on 3D Gaussian Splatting Compression Methods

### v1

@@ Abstract
Method TanksAndTemples MipNeRF360 DeepBlending SyntheticNeRF PSNR SSIM LPIPS Size MB PSNR SSIM LPIPS Size MB PSNR SSIM LPIPS Size MB PSNR SSIM LPIPS Size MB HAC-highrate 24.4 0.853 0.177 11.2 27.77 0.811 0.23 21.9 30.3 0.906 0.258 6.3 33.71 0.968 0.034 1.9 HAC-lowrate 24 0.846 0.187 8.1 27.5 0.807 0.238 15.3 30 0.902 0.269 4.3 33.2 0.967 0.037 1.2 LightGaussian 23.1 0.817 0.231 22 27.3 0.805 0.243 42 32.7 0.965 0.037 7.8 Compact3DGS 23.3 0.831 0.201 39.4 27.1 0.798 0.247 48.8 29.8 0.901 0.258 43.2 33.3 0.968 0.034 5.5 Compact3DGS+PP 23.3 0.831 0.202 20.9 27 0.797 0.247 29.1 29.7 0.9 0.258 23.8 Morgenstern et al. 25.63 0.864 0.208 21.4 27.6 0.814 0.22 40.3 30.4 0.909 0.258 16.8 33.7 0.969 0.031 4.1 Morgenstern et al. w/o SH 25.3 0.857 0.217 8.2 27 0.803 0.232 16.7 30.5 0.908 0.261 5.5 31.8 0.961 0.04 2 Navaneet et al. 16K 23.4 0.836 0.2 12 27 0.804 0.243 18 29.9 0.906 0.252 12 Navaneet et al. 32K 23.4 0.838 0.198 13 27.1 0.806 0.24 19 29.9 0.907 0.251 13 Compressed3D 23.3 0.832 0.194 17.3 27 0.801 0.238 28.8 29.4 0.898 0.253 25.3 32.9 0.967 0.033 3.7 Reduced3DGS 23.6 0.84 0.188 14 27.1 0.809 0.226 29 29.6 0.902 0.249 18

### v2

@@ Abstract
Method
Rank
TanksAndTemples
MipNeRF360
DeepBlending
SyntheticNeRF
PSNR
SSIM
LPIPS
Size
MB
PSNR
SSIM
LPIPS
Size
MB
PSNR
SSIM
LPIPS
Size
MB
PSNR
SSIM
LPIPS
Size
MB
HAC-lowrate
3.0
24.04
0.846
0.187
8.1
27.53
0.807
0.238
15.3
29.98
0.902
0.269
4.3
33.24
0.967
0.037
1.2
HAC-highrate
3.2
24.40
0.853
0.177
11.2
27.77
0.811
0.230
21.9
30.34
0.906
0.258
6.3
33.71
0.968
0.034
1.9
LightGaussian
8.9
23.11
0.817
0.231
22.0
27.28
0.805
0.243
42.0
32.725
0.965
0.037
7.8
EAGLES
8.8
23.37
0.84
0.20
29.0
27.23
0.81
0.24
54.0
29.86
0.91
0.25
52.0
EAGLES-Small
9.9
23.1
0.82
0.22
19.0
26.94
0.80
0.25
47.0
29.92
0.90
0.25
33.0
Compact3DGS
9.2
23.32
0.831
0.201
39.4
27.08
0.798
0.247
48.8
29.79
0.901
0.258
43.2
33.33
0.968
0.034
5.5
Compact3DGS+PP
9.3
23.32
0.831
0.202
20.9
27.03
0.797
0.247
29.1
29.73
0.900
0.258
23.8
Morgenstern et al.
5.1
25.63
0.864
0.208
21.4
27.64
0.814
0.220
40.3
30.35
0.909
0.258
16.8
33.70
0.969
0.031
4.1
Morgenstern et al. w/o SH
4.2
25.27
0.857
0.217
8.2
27.02
0.803
0.232
16.7
30.50
0.908
0.261
5.5
31.75
0.961
0.040
2.0
Navaneet et al. 16K
5.5
23.39
0.836
0.200
12.0
27.03
0.804
0.243
18.0
29.90
0.906
0.252
12.0
Navaneet et al. 32K
5.2
23.44
0.838
0.198
13.0
27.12
0.806
0.240
19.0
29.90
0.907
0.251
13.0
Compressed3D
7.0
23.324
0.832
0.194
17.3
26.981
0.801
0.238
28.8
29.381
0.898
0.253
25.3
32.936
0.967
0.033
3.7
Reduced3DGS
5.9
23.57
0.84
0.188
14.0
27.1
0.809
0.226
29.0
29.63
0.902
0.249
18.0
Note: The best methods in each category is highlighted. The ranks represent the average rankings of the methods across all available datasets. The quality metrics PSNR, SSIM, and LPIPS are equally weighted with the model size, meaning they each contribute one-sixth to the ranks, while the size contributes half.
@@
EAGLES: Efficient Accelerated 3D Gaussians with Lightweight EncodingS
This approach aims to decrease both storage and runtime memory costs while enhancing training and rendering speeds and maintaining view synthesis quality. The authors observed that the color and rotation attributes account for over 80
@@
F-3DGS: Factorized Coordinates and Representations for 3D Gaussian Splatting
The paper introduces a novel 3D Gaussian compression method using structured coordinates and decomposed representations through factorization. Inspired by tensor or matrix factorization techniques, this method generates 3D coordinates via a tensor product of 1D or 2D coordinates, enhancing spatial efficiency. It extends factorization to include attributes like color, scale, rotation, and opacity, which compresses the model size while maintaining essential characteristics. A binary mask is employed to eliminate non-essential Gaussians, significantly accelerating training and rendering speeds. Experiments demonstrate that this approach reduces storage requirements by over 90<br><br> <span style="color:red"><b>Note:</b> This paper is currently not included in the survey table because it shows unusually high results in the Tanks and Temples dataset and reports higher results for the original 3DGS than those in the original publication. This raises the possibility that their testing methods may differ from those used in other papers.</span> [13]